# Position Paper: What Can Large Language Models Tell Us about Time Series Analysis

### v1

@@ Abstract
Time series analysis is essential for comprehending the complexities inherent in various real-world systems and applications. Although large language models (LLMs) have recently made significant strides, the development of artificial general intelligence (AGI) equipped with time series analysis capabilities remains in its nascent phase. Most existing time series models heavily rely on domain knowledge and extensive model tuning, predominantly focusing on prediction tasks. In this paper, we argue that current LLMs have the potential to revolutionize time series analysis, thereby promoting efficient decision-making and advancing towards a more universal form of time series analytical intelligence. Such advancement could unlock a wide range of possibilities, including modality switching and time series question answering. We encourage researchers and practitioners to recognize the potential of LLMs in advancing time series analysis and emphasize the need for trust in these related efforts. Furthermore, we detail the seamless integration of time series analysis with existing LLM technologies and outline promising avenues for future research.
@@ 1 Introduction
The recent emergence of large language models (LLMs), such as Llama (Touvron et al., 2023a, b) and GPT-4 (Achiam et al., 2023), have swept through and propelled advancements in various interdisciplinary fields (Zhao et al., 2023). Their outstanding zero-shot capabilities (Kojima et al., 2022), along with emerging reasoning and planning abilities (Wang et al., 2023a), have garnered increasing attention. However, their focus has primarily been on text sequences. The exploration of extending LLMs’ capabilities to accommodate and process more data modalities, such as images (Zhang et al., 2023b) and graphs (Chen et al., 2023c), has begun to receive preliminary attention.
With the integration of LLMs, time series analysis is undergoing significant transformation (Jin et al., 2023b). Time series models are conventionally designed for specific tasks, depend heavily on prior domain knowledge and extensive model tuning, lacking assurances of effective updates and validations (Zhou et al., 2023a). Conversely, LLMs hold enormous potential not only to improve prediction performance (Jin et al., 2024) but also to support cross-disciplinary (Yan et al., 2023), interactive (Xue et al., 2023), and interpretative (Gu et al., 2023) analyses. By aligning time series and natural language, large language and specialistic time series models constitute a new technology paradigm, where the LLM is prompted with both time series and text-based instructions. In this paradigm, time series and textual information provide essential contexts, LLMs contribute internal knowledge and reasoning capabilities, and pre-trained time series models offer fundamental pattern recognition assurances. This novel integration is depicted in Figure 1, where the successful amalgamation of these components showcases the potential for a general-purpose, unified system in next-generation time series analysis.
@@
Given the remarkable capabilities emerging in recent research (Jin et al., 2023b), we believe that the field of time series analysis research is undergoing an exciting transformative moment. Our standpoint is that LLMs can serve as the central hub for understanding and advancing time series analysis. Specifically, we present key insights that LLMs can profoundly impact time series analysis in three fundamental ways: (1) as effective data and model enhancers, augmenting time series data and existing approaches with enhanced external knowledge and analytical prowess; (2) as superior predictors, utilizing their extensive internal knowledge and emerging reasoning abilities to benefit a range of prediction tasks; and (3) as next-generation agents, transcending conventional roles to actively engage in and transform time series analysis. We advocate attention to related research and efforts, moving towards more universal intelligent systems for general-purpose time series analysis. To this end, we thoroughly examine relevant literature, present and discuss potential formulations of LLM-centric time series analysis to bridge the gap between the two. We also identify and outline prospective research opportunities and challenges, calling for greater commitment and exploration in this promising interdisciplinary field.
@@ 2 Background
This section provides an overview of the fundamental concepts in time series analysis and large language models. Furthermore, it outlines a developmental roadmap for time series analytical models, tracing the progression from traditional statistical methods to advanced, next-generation LLM-centric approaches, thereby synthesizing the foundational principles of both fields.
@@ 2.1 Time Series Analysis
Time series analytics is crucial for deriving insights from data, with recent deep learning advancements spurring a rise in neural network-based methods (Wen et al., 2023). These methods focus on modeling complex inter-temporal and/or inter-variable relationships in time series (Zhang et al., 2023c; Jin et al., 2023b), aiding in tasks like forecasting, classification, anomaly detection, and imputation. Forecasting predicts future values, classification categorizes series by patterns, anomaly detection identifies anomalous events, and imputation estimates missing data. Beyond these tasks, emerging research has shown promise in modality switching and question answering (Xue & Salim, 2023; Jin et al., 2024; Yang et al., 2022a). These novel approaches highlight the potential for cross-disciplinary, interactive, and interpretative advancements in time series analytics. Such advancements open a realm of possibilities in practical applications, such as (zero-shot) medical question answering (Yu et al., 2023a; Oh et al., 2023) and intelligent traffic agents (Da et al., 2023b; Lai et al., 2023).
@@ 2.2 Large Language Models
Large language models exhibit emergent abilities that set them apart from traditional neural networks. These abilities, present in large models but not in smaller ones, are a significant aspect of LLMs (Wei et al., 2022a). Three key emergent abilities of LLMs include: (1) In-context learning (ICL), introduced by GPT-3 (Brown et al., 2020), allowing LLMs to generate relevant outputs for new instances using instructions and examples without additional training; (2) Instruction following, where LLMs, through instruction tuning, excel at novel tasks presented in an instructional format, enhancing their generalization (Sanh et al., 2021); (3) Step-by-step reasoning, where LLMs use strategies like chain-of-thought (CoT) (Wei et al., 2022b) or other prompting strategies (Yao et al., 2023; Besta et al., 2023) to address complex tasks requiring multiple reasoning steps.
@@ 2.3 Research Roadmap
Time series analytical model development spans four generations: (1) statistical models, (2) deep neural networks, (3) pre-trained models, and (4) LLM-centric models, as shown in Figure 2. This categorization hinges on the evolving task-solving capabilities of each model generation. Traditional analytics relied on statistical models like ARIMA (Shumway et al., 2017) and Holt-Winters (Kalekar et al., 2004), optimized for small-scale data and based on heuristics like stationarity and seasonality (Hamilton, 2020). These models assumed past trends would continue into the future. Deep neural networks, like recurrent and temporal convolution neural networks (Gamboa, 2017), processed larger, complex datasets, capturing non-linear and long-term dependencies without heavy reliance on prior knowledge, thus transforming predictive time series analysis. Recent research like TimeCLR (Yeh et al., 2023) introduced pre-training on diverse, large-scale time series data, allowing fine-tuning for specific tasks with relatively smaller data samples (Jin et al., 2023b), reducing the time and resources required for model training. This allows for the application of sophisticated models in scenarios where collecting large-scale time series data is challenging. Despite the successes of previous generations, we posit that the emergence of LLMs is set to revolutionize time series analysis, shifting it from predictive to general intelligence. LLM-centric models, processing both language instructions and time series (Jin et al., 2024; Anonymous, 2024a), extend capabilities to general question answering, interpretable predictions, and complex reasoning, moving beyond conventional predictive analytics.
@@ 3 LLM-assisted Enhancer for Time Series
Numerous time series analysis models have been devised to address temporal data. LLMs, owing to vast internal knowledge and reasoning capabilities, seamlessly enhance both aspects. Hence, we can intuitively distinguish LLM-assisted enhancer methods from data and model perspectives.
@@ 3.1 Data-Based Enhancer
LLM-assisted enhancers not only enhance data interpretability but also provide supplementary improvements, facilitating a more thorough understanding and effective use of time series data. For interpretability, LLMs offer textual descriptions and summaries, helping to understand patterns and anomalies in time series data. Examples include LLM-MPE (Liang et al., 2023) for human mobility data, SignalGPT (Liu et al., 2023a) for biological signals, and Insight Miner (Zhang et al., 2023f) for trend mining. Additionally, AmicroN (Chatterjee et al., 2023) and SST (Ghosh et al., 2023) use LLMs for detailed sensor and spatial time series analysis. Supplementary enhancements involve integrating diverse data sources, enriching time series data context and improving model robustness, as explored in (Yu et al., 2023b) and (Fatouros et al., 2024) for financial decision-making. Such enhancements help improve domain models’ inherent capabilities and make them more robust.
@@ 3.2 Model-Based Enhancer
Model-based enhancers aim to augment time series models by addressing their limitations in external knowledge and domain-specific contexts. Transferring knowledge from LLMs boosts the performance of domain models in handling complex tasks. Such approaches often employ a dual-tower model, like those in (Qiu et al., 2023b; Li et al., 2023b; Yu et al., 2023a), use frozen LLMs for electrocardiogram (ECG) analysis. Some methods further utilize contrastive learning to achieve certain alignments. For example, IMU2CLIP (Moon et al., 2023) aligns text and video with sensor data, while STLLM (Anonymous, 2024b) enhances spatial time series prediction. Another line of work utilizes prompting techniques to harness the inferential decision-making capability of LLMs. For instance, TrafficGPT (Zhang et al., 2023e) exemplifies decision analysis, integrating traffic models with LLMs for user-tailored solutions, offering detailed insights to enhance system interpretability.
@@ 3.3 Discussion
LLM-assisted enhancers effectively address the inherent sparsity and noise characteristics of time series data, providing existing time series models with more effective external knowledge and analytical capabilities. Moreover, this technology is plug-and-play, enabling flexible assistance for real-world time series data and model challenges. However, a notable hurdle is that using LLM as an enhancer introduces significant time and cost overheads when dealing with large-scale datasets. In addition, the inherent diversity and range of application scenarios in time series data add layers of complexity to the creation of universally effective LLM-assisted enhancers.
@@ 4.1 Tuning-Based Predictor
Tuning-based predictors use accessible LLM parameters, typically involving patching and tokenizing numerical signals and related text data, followed by fine-tuning for time series tasks. Figure 3(a) shows this process: (1) with a $\operatorname{Patching}(\cdot)$ operation (Nie et al., 2022), a time series is chunked to form patch-based tokens $\mathcal{X}_{inp}$. An additional option is to perform $\operatorname{Tokenizer}(\cdot)$ operation on time series-related text data to form text sequence tokens $\mathcal{T}_{inp}$; (2) time series patches (and optional text tokens) are fed into the LLM with accessible parameters; (3) an extra task layer, denoted as $Task(\cdot)$, is finally introduced to perform different analysis tasks with the instruction prompt $P$. This process is formulated below:
@@
Adapting pre-trained large language models directly to raw time series numerical signals for downstream time series analysis tasks is often counterintuitive due to the inherent modality gap between text and time series data. Nevertheless, FPT (Zhou et al., 2023a) and similar studies found that LLMs, even when frozen, can perform comparably in time series tasks due to the self-attention mechanism’s universality. Others, like GATGPT (Chen et al., 2023b) and ST-LLM (Liu et al., 2024a), applied these findings to spatial-temporal data, while UniTime (Liu et al., 2024b) used manual instructions for domain identification. This allows them to handle time series data with different characteristics and distinguish between different domains.
However, the above methods all require modifications that disrupt the parameters of the original LLMs, potentially leading to catastrophic forgetting. In contrast, another line of work, inspired by this, aims to avoid this by introducing additional lightweight adaptation layers. Time-LLM (Jin et al., 2024) uses text data as a prompt prefix and reprograms input time series into language space, enhancing LLM’s performance in various forecasting scenarios. TEST (Sun et al., 2024) tackles inconsistent embedding spaces by constructing an encoder for time series data, employing alignment contrasts and soft prompts for efficient fine-tuning with frozen LLMs. LLM4TS (Chang et al., 2023) integrates multi-scale time series data into LLMs using a two-level aggregation strategy, improving their interpretation of temporal information. TEMPO (Cao et al., 2023) combines seasonal and trend decompositions with frozen LLMs, using prompt pooling to address distribution changes in forecasting non-stationary time series.
@@ 4.2 Non-Tuning-Based Predictor
Non-tuning-based predictors, suitable for closed-source models, involve preprocessing time series data to fit LLM input spaces. As Figure  3(b) illustrates, this typically involves two steps: (1) preprocessing raw time series, including optional operations such as prompt $\operatorname{Template}(\cdot)$ and customized $\operatorname{Tokenizer}(\cdot)$; (2) feeding the processed inputs $\mathcal{X}_{inp}$ into the LLM to obtain responses. A $\operatorname{Parse}(\cdot)$ function is then employed to retrieve prediction labels. This process is formulated below:
@@
(Spathis & Kawsar, 2023) initially noted that LLM tokenizers, not designed for numerical values, separate continuous values and ignore their temporal relationships. They suggested using lightweight embedding layers and prompt engineering as solutions. Following this, LLMTime (Gruver et al., 2023) introduced a novel tokenization approach, converting tokens into flexible continuous values, enabling non-tuned LLMs to match or exceed zero-shot prediction performance in domain-specific models. This success is attributed to LLMs’ ability to represent multimodal distributions. Using in-context learning, evaluations were performed in tasks like sequence transformation and completion. (Mirchandani et al., 2023) suggested that LLMs’ capacity to handle abstract patterns positions them as foundational general pattern machines. This has led to applying LLMs in areas like human mobility mining (Wang et al., 2023c; Zhang et al., 2023g), financial forecasting (Lopez-Lira & Tang, 2023), and health prediction (Kim et al., 2024).

### v2

# Position: What Can Large Language Models Tell Us about Time Series Analysis

@@ Abstract
Time series analysis is essential for comprehending the complexities inherent in various real-world systems and applications. Although large language models (LLMs) have recently made significant strides, the development of artificial general intelligence (AGI) equipped with time series analysis capabilities remains in its nascent phase. Most existing time series models heavily rely on domain knowledge and extensive model tuning, predominantly focusing on prediction tasks. In this paper, we argue that current LLMs have the potential to revolutionize time series analysis, thereby promoting efficient decision-making and advancing towards a more universal form of time series analytical intelligence. Such advancement could unlock a wide range of possibilities, including time series modality switching and question answering. We encourage researchers and practitioners to recognize the potential of LLMs in advancing time series analysis and emphasize the need for trust in these related efforts. Furthermore, we detail the seamless integration of time series analysis with existing LLM technologies and outline promising avenues for future research.
@@ 1 Introduction
The recent emergence of large language models (LLMs), such as Llama (Touvron et al., 2023a, b) and GPT-4 (Achiam et al., 2023), have swept through and propelled advancements in various interdisciplinary fields (Zhao et al., 2023). Their outstanding zero-shot capabilities (Kojima et al., 2022), along with emerging reasoning and planning abilities (Wang et al., 2024), have garnered increasing attention. However, their focus has primarily been on text sequences. The exploration of extending LLMs’ capabilities to accommodate and process more data modalities, such as images (Zhang et al., 2024a) and graphs (Chen et al., 2024), has begun to receive preliminary attention.
With the integration of LLMs, time series analysis is undergoing significant transformation (Jin et al., 2023b; Liang et al., 2024). Time series models are conventionally designed for specific tasks, depend heavily on prior domain knowledge and extensive model tuning, lacking assurances of effective updates and validations (Zhou et al., 2023a). Conversely, LLMs hold enormous potential not only to improve prediction performance (Jin et al., 2024) but also to support cross-disciplinary (Yan et al., 2023), interactive (Xue et al., 2023), and interpretative (Gu et al., 2024) analyses. By aligning time series and natural language, large language and foundation/specialistic time series models constitute a new technology paradigm, where the LLM is prompted with both time series and textual instructions. In this paradigm, time series and textual information provide essential contexts, LLMs contribute internal knowledge and reasoning capabilities, and pre-trained time series models offer fundamental pattern recognition assurances. This novel integration is depicted in Figure 1, where the successful amalgamation of these components showcases the potential for a general-purpose, unified system in next-generation time series analysis.
@@
Given the remarkable capabilities emerging in recent research (Jin et al., 2023b; Liang et al., 2024; Zhang et al., 2024e), we believe that the field of time series analysis research is undergoing an exciting transformative moment. Our standpoint is that LLMs can act as the central hub for understanding and advancing the analysis of time series data. Specifically, we present key insights that LLMs can profoundly impact time series analysis in three fundamental ways with their capability boundaries illustrated in Figure 2: (1) as effective data and model enhancers, augmenting time series data and existing approaches with enhanced external knowledge and analytical prowess; (2) as superior predictors, utilizing their extensive internal knowledge and emerging reasoning abilities to benefit a range of prediction tasks; and (3) as next-generation agents, transcending conventional roles to actively engage in and transform time series analysis. We advocate attention to related research and efforts, moving towards more universal intelligent systems for general-purpose time series analysis. To this end, we thoroughly examine relevant literature, present and discuss potential formulations of LLM-centric time series analysis to bridge the gap between the two. We also identify and outline prospective research opportunities and challenges, calling for greater commitment and exploration in this promising interdisciplinary field.
Figure 2: Task-solving capability boundaries on the roles of LLMs for time series analysis: as data/model enhancers, effective predictors, or next-generation agents.
@@ 2.1 Time Series Analysis
Time series analysis is crucial for deriving insights from data, with recent deep learning advancements spurring a rise in neural network-based methods (Wen et al., 2023). These methods focus on modeling complex inter-temporal and/or inter-variable relationships in time series (Zhang et al., 2024b; Jin et al., 2023b), aiding in tasks like forecasting, classification, anomaly detection, and imputation. Forecasting predicts future values, classification categorizes series by patterns, anomaly detection identifies anomalous events, and imputation estimates missing data. Emerging research has also shown promise in time series modality switching and question answering (Xue & Salim, 2023; Jin et al., 2024; Yang et al., 2022a). These novel approaches highlight the potential for cross-disciplinary, interactive, and interpretative advancements in time series analytics. Such advancements open a realm of possibilities in practical applications, such as (zero-shot) medical question answering (Yu et al., 2023a; Oh et al., 2024) and intelligent traffic agents (Da et al., 2024; Lai et al., 2023).
@@ 2.2 Large Language Models
Large language models exhibit emergent abilities that set them apart from traditional neural networks. These abilities, present in large models but not in smaller ones, are a significant aspect of LLMs (Wei et al., 2022a). Three key emergent abilities of LLMs include: (1) in-context learning (ICL), introduced by GPT-3 (Brown et al., 2020), allowing LLMs to generate relevant outputs for new instances using instructions and examples without additional training; (2) instruction following, where LLMs, through instruction tuning, excel at novel tasks presented in an instructional format, enhancing their generalization (Victor et al., 2022); (3) step-by-step reasoning, where LLMs use strategies like chain-of-thought (CoT) (Wei et al., 2022b) or other prompting strategies (Yao et al., 2024; Besta et al., 2023) to address complex tasks requiring multiple reasoning steps.
@@ 2.3 Research Roadmap
Time series analytical model development spans four generations: (1) statistical models, (2) deep neural networks, (3) pre-trained models, and (4) LLM-centric models, as shown in Figure 3. This categorization hinges on the evolving task-solving capabilities of each model generation. Traditional analytics relied on statistical models like ARIMA (Shumway et al., 2017) and Holt-Winters (Kalekar et al., 2004), optimized for small-scale data and based on heuristics like stationarity and seasonality (Hamilton, 2020). These models assumed past trends would continue into the future. Deep neural networks, like recurrent and temporal convolution neural networks (Gamboa, 2017), processed larger, complex datasets, capturing non-linear and long-term dependencies without heavy reliance on prior knowledge, thus transforming predictive time series analysis. Recent research like TimeCLR (Yeh et al., 2023) introduced pre-training on diverse, large-scale time series data, allowing fine-tuning for specific tasks with relatively smaller data samples (Jin et al., 2023b), reducing the time and resources required for model training. This allows for the application of sophisticated models in scenarios where collecting large-scale time series data is challenging. Despite the successes of previous generations, we posit that the emergence of LLMs is set to revolutionize time series analysis, shifting it from predictive to general intelligence. LLM-centric models, processing both language instructions and time series (Jin et al., 2024; Cheng & Chin, 2024), extend capabilities to general question answering, interpretable predictions, and complex reasoning, moving beyond conventional predictive analytics.
@@ 3 LLM-assisted Enhancer for Time Series
The role of LLM-assisted enhancers delves into whether LLMs can augment our understanding of time series data and extend the knowledge of existing models. Numerous methods have been devised to address temporal data, but the vast internal knowledge and reasoning capabilities of LLMs may significantly enhance both data understanding and model performance; thus, we intuitively distinguish LLM-assisted enhancers from data and model perspectives.
@@ 3.1 Data-based Enhancer
LLM-assisted enhancers not only enhance data interpretability but also provide supplementary improvements, facilitating a more thorough understanding and effective use of time series data. For interpretability, LLMs offer textual descriptions and summaries, helping to understand patterns and anomalies in time series data. Examples include LLM-MPE (Liang et al., 2023) for human mobility data, SignalGPT (Liu et al., 2023a) for biological signals, and Insight Miner (Zhang et al., 2023c) for trend mining. Additionally, AmicroN (Chatterjee et al., 2023) and SST (Ghosh et al., 2023) use LLMs for detailed sensor and spatial time series analysis. Supplementary enhancements involve integrating diverse data sources, enriching time series data context and improving model robustness, as explored in (Yu et al., 2023b) and (Fatouros et al., 2024) for financial decision-making. Such enhancements help improve domain models’ inherent capabilities and make them more robust.
@@ 3.2 Model-based Enhancer
Model-based enhancers aim to augment time series models by addressing their limitations in external knowledge and domain-specific contexts. Transferring knowledge from LLMs boosts the performance of domain models in handling complex tasks. Such approaches often employ a dual-tower model, like those in (Qiu et al., 2023b; Li et al., 2024), use frozen LLMs for electrocardiogram (ECG) analysis. Some methods further utilize contrastive learning to achieve certain alignments. For example, IMU2CLIP (Moon et al., 2023) aligns text and video with sensor data, while STLLM (Zhang et al., 2024c) enhances spatial time series prediction. Another branch utilizes prompting techniques to harness the inferential decision-making capability of LLMs. For instance, TrafficGPT (Zhang et al., 2024d) exemplifies decision analysis, integrating traffic models with LLMs for user-tailored solutions, offering detailed insights to enhance system interpretability.
@@ 3.3 Discussion
LLM-assisted enhancers effectively address the inherent sparsity and noise characteristics of time series data while also providing existing time series models with enhanced external knowledge and analytical capabilities. Time series, perhaps more than other modalities, benefits significantly from data enhancements due to its potentially lower information density. For example, comparing raw data volumes, video data inherently contain vastly more information per second than typical time series data such as audio. This disparity underscores the importance of time series data augmentation in many use cases, where LLMs can leverage both internal and external knowledge, along with their inherent reasoning capabilities, to address this challenge. On the other hand, the development of foundation models for time series is still in its infancy, presenting an opportunity for LLMs to extend the knowledge boundaries of existing time series models. Moreover, this technology is plug-and-play, enabling flexible assistance for real-world time series data and model challenges. However, a notable hurdle is that using LLM as an enhancer introduces significant time and cost overheads when dealing with large-scale datasets. In addition, the inherent diversity and range of application scenarios in time series data add layers of complexity to the creation of universally effective LLM-assisted enhancers.
@@ 4.1 Tuning-based Predictor
Tuning-based predictors use accessible LLM parameters, typically involving patching and tokenizing numerical signals and related text data, followed by fine-tuning for time series tasks. Figure 4(a) shows this process: (1) with a $\operatorname{Patching}(\cdot)$ operation (Nie et al., 2022), a time series is chunked to form patch-based tokens $\mathcal{X}_{inp}$. An additional option is to perform $\operatorname{Tokenizer}(\cdot)$ operation on time series-related text data to form text sequence tokens $\mathcal{T}_{inp}$; (2) time series patches (and optional text tokens) are fed into the LLM with accessible parameters; (3) an extra task layer, denoted as $Task(\cdot)$, is finally introduced to perform different analysis tasks with the instruction prompt $P$ (not shown in the figure). This process is formulated below:
@@
Adapting out-of-box LLMs directly to raw time series numerical signals for downstream time series analysis tasks is often counterintuitive due to the inherent modality gap between text and time series data. Nevertheless, OFA (Zhou et al., 2023a) and similar studies found that LLMs, even when frozen, can perform comparably in time series tasks due to the self-attention mechanism’s universality. Others, like GATGPT (Chen et al., 2023b) and ST-LLM (Liu et al., 2024a), applied these findings to spatial-temporal data, while UniTime (Liu et al., 2024b) used manual instructions for domain identification. This allows them to handle time series data with different characteristics and distinguish between different domains.
However, the above methods all require modifications that disrupt the parameters of the original LLMs, potentially leading to catastrophic forgetting. In contrast, another line of work, inspired by this, aims to avoid this by introducing additional lightweight adaptation layers. Time-LLM (Jin et al., 2024) uses text data as a prompt prefix and reprograms input time series into language space, enhancing LLM’s performance in various forecasting scenarios. TEST (Sun et al., 2024) tackles inconsistent embedding spaces by constructing an encoder for time series data, employing alignment contrasts and soft prompts for efficient fine-tuning with frozen LLMs. TEMPO (Cao et al., 2023) combines seasonal and trend decompositions with frozen LLMs, using prompt pooling to address distribution changes in forecasting non-stationary time series.
Figure 4: Categories of LLM-centered predictor.
@@ 4.2 Non-tuning-based Predictor
Non-tuning-based predictors, suitable for closed-source models, involve preprocessing time series data to fit the input spaces of LLMs. As Figure  4(b) illustrates, this typically involves two steps: (1) preprocessing raw time series, including optional operations such as prompt $\operatorname{Template}(\cdot)$ and customized $\operatorname{Tokenizer}(\cdot)$; (2) feeding the processed inputs $\mathcal{X}_{inp}$ into the LLM to obtain responses. A $\operatorname{Parse}(\cdot)$ function is then employed to retrieve prediction labels. This process is formulated below:
@@
(Spathis & Kawsar, 2023) initially noted that LLM tokenizers, not designed for numerical values, separate continuous values and ignore their temporal relationships. They suggested using lightweight embedding layers and prompt engineering as solutions. Following this, LLMTime (Gruver et al., 2023) introduced a novel tokenization approach, converting tokens into flexible continuous values, enabling non-tuned LLMs to match or exceed zero-shot prediction performance in domain-specific models. This success is attributed to LLMs’ ability to represent multimodal distributions. Using in-context learning, evaluations were performed in tasks like sequence transformation and completion. (Mirchandani et al., 2023) suggested that LLMs’ capacity to handle abstract patterns positions them as foundational general pattern machines. This has led to applying LLMs in areas like human mobility mining (Wang et al., 2023b; Zhang et al., 2023d), financial forecasting (Lopez-Lira & Tang, 2023), and health prediction (Kim et al., 2024).